OpenStudy (anonymous):

The function of t(x) = x + 6 determine how many cans of soup a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew used c(t(x)) to find the amount of money to spend on soup. THe function c(x) = 2x +4. Solve for how much money must be spent when the crew is going to work 4 shifts

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

t(x) = x+ 6 As crew worked for 4 shifts then total time will be t(4) = 4+6 t(4) = 10 So then total cost will be c(t(4)) = 2(10) + 4 c(t(4)) = 24